OpenStudy (anonymous):

solve the system by the elimination method 4x-9y=18.5 7y-2x=10.5

OpenStudy (anonymous):

x=22.4 y=7.9

OpenStudy (anonymous):

no one is presenting these correctly, they have to be in a paired order of fractions

OpenStudy (anonymous):

y=13,16

OpenStudy (anonymous):

Let 1. => 4x - 9y = 18.5 and 2. => 7y-2x = 10.5 2 X 2. => 14y -4x = 21 Then 1. + 2 X 2. Therefore, (4x-4x) + (-9y + 14y) = 39.5 0x + 5y = 39.5 y = 39.5/5 = 79/10

OpenStudy (anonymous):

y=5,26

OpenStudy (anonymous):

Then such solve for x using y value.

OpenStudy (anonymous):

no they were not

OpenStudy (anonymous):

Start off by multiplying equation 2 by 2. This will cancel the x's. We now have: 4x-9y = 18.5 -4x + 14y = 21 If we add together, we get: 5y = 39.5 Dividing through, we have: (39/2)/5 = y To simplify the fraction, we just multiply 39/2 * 1/5 Which is 39/10. Now plug this value for y into either equation above to get x. 7(39/10)-2x = 10.5 273/10-2x=105/10 - By making a common denominator. -2x = -84/5 By subtracting 273/10 from both sides. DIviding both sides by -2 gives us: x = (-84/5)/-2 Now by multiplying -84/5 * -1/2 we have: x = -84/10 Bringing to LCD, we have: x = -42/5 If this helps please give me a medal. http://www.tutorsean.net for free math videos.

OpenStudy (anonymous):

There is your METAL bro Cali23760 is mean :(.